OpenStudy (anonymous):

How do you write the equation f(x)=x62+6x+3 in the form f(x)=a(x-h)^2 + k?

OpenStudy (amoodarya):

do you mean 6x^2+6x+3? or x^62+6x+3 ?

hartnn (hartnn):

\( f(x)=x^2+6x+3 \)

OpenStudy (anonymous):

f(x)= x^2+6x+3

hartnn (hartnn):

do you know how to complete the square ?

OpenStudy (anonymous):

one half b squared?

hartnn (hartnn):

yes! here, b=.... ?

OpenStudy (anonymous):

b=6

hartnn (hartnn):

and (b/2)^2 =... ?

OpenStudy (anonymous):

9

hartnn (hartnn):

so, add and subtract 9 in the function, f(x)= x^2+6x+3

hartnn (hartnn):

f(x)= x^2+6x+9+(-9+3) got this ?

OpenStudy (anonymous):

isn't it in the other side of the equation?

hartnn (hartnn):

'it' what ?

OpenStudy (anonymous):

the 9-3

hartnn (hartnn):

lets keep f(x) on left and other things on right only...

OpenStudy (anonymous):

ok

hartnn (hartnn):

so, notice that x^2+6x+9 is a perfect square... of ?

hartnn (hartnn):

\(f(x)=x^2+6x+9 +(-6) \\ =(....)^2 + (-6)\) can u fill the blank ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

(x+3)

hartnn (hartnn):

yes, so you got that form you required.... \(f(x)=(x+3)^2-6 \\ = 1.(x- (-3))^2+ (-6)\)

OpenStudy (anonymous):

Jun has 1200 feet of fencing materials to build a pen for his sheep and goats. The pen will be a rectangle with a divider down the middle. What should be the maximum are for the pen?

OpenStudy (anonymous):

heeeeeeeeellllllllllllppppppppppppp

hartnn (hartnn):

is that calculus related ? do you know how to differentiate ?

OpenStudy (anonymous):

i'm not that really good in math..........

hartnn (hartnn):

and i would like to request you to make a new post for new question from next time..

hartnn (hartnn):

i meant what is the topic for that question ?

OpenStudy (anonymous):

to know its maximum area

hartnn (hartnn):

that can be found using several methods...are u ever taught derivatives ?

OpenStudy (anonymous):

yeah we have discussed that

OpenStudy (anonymous):

but i'm so confused using derivatives in its application

hartnn (hartnn):

good, do you know derivative of x^2 and x ?

OpenStudy (anonymous):

yeah i know simple derivatives

hartnn (hartnn):

ok, let go to your new post..